the perimeter of a rectangle is 146m the length is 4m more than 2 times the width. find the dimensions
Answer by checkley77(12844)

L=2W+4
P=2L+2W
146=2(2W+4)+2W
146=4W+8+2W
146=6W+8
6W=146-8
6W=138
W=136/6
W=23 ANS. FOR THE WIDTH.
L=2*23+4=46+4=50 ANS. FOR THE LENGTH.
PROOF:
146=2*50+2*23
146=100+46
146=146